# Multiplication of 0-1 matrices via clustering

### v1

@@ Abstract
We study applications of clustering (in particular the $k$-center clustering problem) in the design of efficient and practical deterministic algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices $A$ and $B$ with clustered rows or columns, respectively. Let $\lambda_{A}$ and $\lambda_{B}$ denote the minimum maximum radius of a cluster in an $\ell$-center clustering of the rows of $A$ and in a $k$-center clustering of the columns of $B,$ respectively. In particular, assuming that the matrices have size $n\times n$, we obtain the following results.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within the additive error of at most $2\lambda_{A}$ in $O(n^{2}\ell)$ time or at most $2\lambda_{B}$ in $O(n^{2}k)$ time.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(n^{2}\ell)$ time or $O(n^{2}k)$ time such that a query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time or $O(\lambda_{B})$ time, respectively.
@@ 1 Introduction
The arithmetic matrix product of two 0-1 matrices is closely related to the Boolean one of the corresponding Boolean matrices. For square $n\times n$ matrices, both can be computed in $O(n^{2.372})$ time [20]. Both are basic tools in science and engineering. Unfortunately, no truly subcubic practical algorithms for any of them are known. Therefore, many researchers studied the complexity of these products for special input matrices, e.g., sparse or clustered matrices [1, 8, 12, 18, 21], providing faster and often more practical algorithms.
The method of multiplying matrices with clustered rows or columns, proposed for Boolean matrix product in [1] and subsequently generalized in [8, 12], relies on the construction of an approximate spanning tree of the rows of the first input matrix or the columns of the second input matrix in a Hamming space. Then, each column of the product matrix or each its row is computed with the help of a traversal of the tree in time proportional to the total Hamming cost of the tree up to a logarithmic factor. Simply, the next entry in a column or a row in the product matrix can be obtained from the previous one in time roughly proportional to the Hamming distance between the corresponding rows or columns of the first or the second input matrix, respectively. Thus, in case the entire tree cost is substantially subquadratic in $n,$ the total running time of this method becomes substantially subcubic provided that a good approximation of a minimum spanning tree of the rows of the first input matrix or the columns of the second one can be constructed in a substantially subcubic time. As for simplicity and practicality, a weak point of this method is that in order to construct such an approximation relatively quickly, it employs a randomized dimension reduction.
@@
Cohen and Lewis [5] and Drineas et al. [6] used random sampling to approximate arithmetic matrix product. Their papers provide an approximation $D$ of the matrix product $AB$ of two $n\times n$ matrices $A$ and $B$ such that $||AB-D||_{F}=O(||AB||_{F}/\sqrt{c})$, for a parameter $c>1$ (see also [18]). The approximation algorithm in [6] runs in $O(n^{2}c)$ time. Drineas et al. [6] also derived bounds on the entrywise differences between the exact matrix product and its approximation. Unfortunately, the best of these bounds is $\Omega(M^{2}n/\sqrt{c})$, where $M$ is the maximum value of an entry in $A$ and $B.$ Sarlós [19] by using a sketch technique obtained the same Frobenius norm guarantees, also in $O(n^{2}c)$ time. However, he derived stronger individual upper bounds on the additive error of each entry $D_{ij}$ of the approximation matrix $D$. They are of the form $O(||A_{i*}||_{2}||B_{*j}||_{2}/\sqrt{c})$, where $A_{i*}$ and $B_{*j}$ stands for the $i$-th row of $A$ and the $j$-th column of $B,$ respectively, that hold with high probability. More recently, Pagh [18] presented a randomized approximation $\tilde{O}(n(n+c))$-time algorithm for the arithmetic product of $n\times n$ matrices $A$ and $B$ such that each entry of the approximate matrix product differs at most by $||AB||_{F}/\sqrt{c}$ from the correct one. His algorithm first compresses the matrix product to a product of two polynomials and then uses the fast Fourier transform to multiply the polynomials. Subsequently, Kutzkov [17] developed analogous deterministic algorithms employing different techniques. For approximation results related to sparse arithmetic matrix products, see [16, 18].
@@ 1.1 Our contributions
The $k$-center clustering problem in a Hamming space $\{0,1\}^{d}$ is for a set $P$ of $n$ points in $\{0,1\}^{d}$ to find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u),$ where $\mathrm{ham}(v,u)$ stands for the Hamming distace between $v$ and $u,$ (i.e., the number of coordinate positions they differ from each other). Each center in $T$ induces a cluster consisting of all points in $P$ for which it is the nearest center.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within the additive error of at most $2\lambda_{A}$ in $O(pq\ell+pr)$ time if $p\geq r$ or at most $2\lambda_{B}$ in $O(qrk+pr)$ time if $p\leq r$.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(pq\ell+pr)$ time if $p\geq r$ or $O(qrk+pr)$ time if $p\leq r$ such that a query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time if $p\geq r$ or $O(\lambda_{B})$ time if $p\leq r.$
@@ 1.2 Techniques
All our main results rely on the classical, simple $2$-approximation algorithm for the $k$-center clustering problem (furthest neighbor traversal) due to Gonzalez [13] (see also Fact 1). Two of them rely also on the idea of updating the inner product of two vectors $a$ and $b$ in $\{0,1\}^{q}$ over the Boolean or an arithmetic semi-ring to that of two vectors $a^{\prime}$ and $b^{\prime}$ in $\{0,1\}^{q}$, where $a=a^{\prime}$ or $b=b^{\prime},$ in time roughly proportional to $\mathrm{ham}(a,a^{\prime})+\mathrm{ham}(b,b^{\prime})$. The idea has been used in [1, 8, 12]. As in the aforementioned papers, we combine it with a traversal of an approximate minimum spanning tree of the rows or columns of an input matrix in the Hamming space $\{0,1\}^{q}$, where $q$ is the length of the rows or columns (see also Lemma 6).
@@ 1.3 Paper organization
The next section contains basic definitions. Section 3 presents our simple approximation algorithm for the arithmetic product of two 0-1 matrices and the simple preprocessing enabling efficient answers to queries asking for the value of an arbitrary entry of the arithmetic product matrix. Section 4 is devoted to our simple algorithm for the exact arithmetic matrix product of two 0-1 matrices. We conclude with a short discussion on possible extensions of our results.
@@ 2 Preliminaries
The minimum-diameter $k$-clustering problem in a Hamming space $\{0,1\}^{d}$ is as follows: given a set $P$ of $n$ points in $\{0,1\}^{d}$, find a partition of $P$ into $k$ subsets $P_{1},P_{2},\dots,P_{k}$ that minimize $\max_{i\in[k]}\max_{v,u\in P_{i}}\mathrm{ham}(v,u).$ Note that the $k$-center clustering problem could be also termed as the minimum-radius $k$-clustering problem. It is known to be NP-hard and even NP-hard to approximate within $2-\epsilon$ [10, 13].
@@ 3 An approximate arithmetic matrix product of 0-1 matrices
procedure $APPROXMMCLUS(A,B,\ell,k)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, and two positive integers $\ell$ and $k,$ not exceeding $p$ and $r,$ respectively.
@@
  2. 2.
    If $p<r$ then
    1. (a)
      Determine an approximate $k$-center clustering of the columns of the matrix $B$ in $\{0,1\}^{q}.$ For each column $B_{*j}$ of $B,$ set $cen_{k}(B_{*j})$ to the center of the cluster in the $k$-clustering to which $B_{*j}$ belongs.
    2. (b)
      Form the $q\times k$ matrix $B^{\prime},$ where the $j^{\prime}$-th column is the $j^{\prime}$-th center in the approximate $k$-center clustering of the columns of $B.$
    3. (c)
      Compute the arithmetic $p\times k$ matrix product $C^{\prime}$ of $A$ and $B^{\prime}.$
    4. (d)
      For $1\leq i\leq p$ and $1\leq j\leq r,$ set $D_{ij}$ to $C^{\prime}_{ij^{\prime}}$, where the $j^{\prime}$-th column $B^{\prime}_{*j^{\prime}}$ of $B^{\prime}$ is $cen_{k}(B_{*j})$.
@@
Suppose a $2$-approximation algorithm for the $\ell$-center clustering or the $k$-center clustering is used in $APPROXMMCLUS(A,B,\ell,k)$ and $C$ stands for the arithmetic product of $A$ and $B.$ Then, for $1\leq i\leq p$ and $1\leq j\leq r,$ $|C_{ij}-D_{ij}|\leq 2\lambda(A,\ell,row)$ if $p\geq r$ or $|C_{ij}-D_{ij}|\leq 2\lambda(B,k,col)$ if $p<r$.
@@
Suppose $p\geq r.$ For $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is the inner product of $cen_{\ell}(A_{i*})$, where $\mathrm{ham}(A_{i*},cen_{\ell}(A_{i*}))\leq 2\lambda(A,\ell,row)$, with $B_{*j}.$ Hence, $C_{ij}$, which is the inner product of $A_{i*}$ with $B_{*j},$ can differ at most by $2\lambda(A,\ell,row)$ from $D_{ij}.$ The case $p<r$ follows by symmetry. ∎
@@
$APPROXMMCLUS(A,B,\ell,k)$ can be implemented in $O(pq\ell+pr+T(\ell,q,r))$ time if $p\geq r$ or $O(rqk+pr+T(p,q,k))$ time if $p<r.$
@@
Suppose $p\geq r.$ Step 1(a), inclusive the assignment of the closest center to each row of $A$, can be done in $O(pq\ell)$ time by using Fact 1, i.e., the classic algorithm of Gonzalez [13]. Step 1(b) takes $O(\ell q)$ time, which is $O(T(\ell,q,r))$ time. Finally, Step 1(c) takes $T(\ell,q,r)$ time while Step 1(d) can be done in $O(pr)$ time. Thus, the overall time is $O(pq\ell+pr+T(\ell,q,r)).$ The upper time bound $O(rqk+pr+T(p,q,k))$ for the $p<r$ case follows by symmetry. ∎
We can use the straightforward $O(sqt)$-time algorithm for the multiplication of two matrices of sizes $s\times q$ and $q\times t,$ respectively. By $\ell qr\leq pq\ell$ if $p\geq r,$ and $pqk\leq qrk$ if $p<r,$ Lemmata 1 and 2 yield the first part (1) of our first main result and its second part (2) for $p<r.$ The whole second part for $p\leq r$ follows from the first part by $(AB)^{\top}=B^{\top}A^{\top}.$
@@
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. There is a simple deterministic algorithm which provides an approximation of all entries of the arithmetic matrix product of $A$ and $B$ within the additive error of at most:
@@
We slightly extend $APPROXMMCLUS(A,B,\ell,k)$ in order to obtain a preprocessing for answering queries about single entries of the arithmetic matrix product of $A$ and $B.$
procedure $PREPROCMMCLUS(A,B,\ell,k)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, and two positive integers $\ell$ and $k,$ not exceeding $p$ and $r,$ respectively.
Output: The $p\times r$ matrix $D$ returned by $APPROXMMCLUS(A,B,\ell,k)$, and for $1\leq i\leq p$ and $1\leq j\leq r,$ the sets of coordinate indices $ind(A,i)$ or $ind(B,j)$ on which $A_{i*}$ or $B_{*j}$ differ from their cluster centers, respectively.
@@
    If $p\geq r$ then for $1\leq i\leq p,$ determine the set $ind(A,i)$ of coordinate indices on which $A_{i*}$ differs from $cen_{\ell}(A_{i*})$.
  3. 3.
    If $p<r$ then for $1\leq j\leq r,$ determine the set $ind(B,j)$ of coordinate indices on which $B_{*j}$ differs from $cen_{k}(B_{*j}).$
@@
$PREPROCMMCLUS(A,B,\ell,k)$ can be implemented in $O(pq\ell++pr+T(\ell,q,r))$ time if $p>r$ or $O(qrk+pr+T(p,q,k))$ time if $p<r.$
@@
Step 1 can be done in $O(pq\ell+pr+T(\ell,q,r))$ time if $p\geq r$ or $O(rqk+pr+T(p,q,k))$ time otherwise by Lemma 2. Steps 2 and 3 can be easily implemented in $O(pq)$ time and $O(rq)$ time, if $p\geq r$ or $p<r$, respectively. ∎
@@
procedure $QUERYMMCLUS(A,B,\ell,k,i,j)$
Input: The preprocessing done by $PREPROCMMCLUS(A,B,\ell,k)$ for 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, $\ell\in[p]$, $k\in[r],$ and two query indices $i\in[p]$ and $j\in[r].$
@@
  3. 3.
    If $p<r$ then for $m\in ind(B,j)$ do
    1. (a)
      If the $m$-th coordinate of the center assigned to $B_{*j}$ is 00 and $A_{im}=1$ then $C_{ij}\leftarrow C_{ij}+1$.
    2. (b)
      If the $m$-th coordinate of the center assigned to $B_{*j}$ is $1$ and $A_{im}$ is also $1$ then $C_{ij}\leftarrow C_{ij}-1$.
@@
$QUERYMMCLUS(A,B,\ell,k,i,j)$ is correct, i.e., the final value of $C_{ij}$ is the inner product of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
@@
Suppose $p\geq r.$ $C_{ij}$ is initially set to $D_{ij}$, which is the inner product of the center assigned to $A_{i*}$ and $B_{*j}.$ Then, $C_{ij}$ is appropriately corrected by increasing or decreasing with $1$ for each coordinate index $m\in ind(A,i)$ which contributes $1$ to the inner product of $A_{i*}$ and $B_{*j}$ and 00 to the inner product of the center of $A_{i*}$ and $B_{*j}$ or vice versa. The correctness of the $p<r$ case follows by symmetric arguments. ∎
@@
$QUERYMMCLUS(A,B,\ell,k,i,j)$ takes $O(\lambda(A,\ell,row))$ time if $p\geq r$ or $O(\lambda(B,k,row))$ time if $p<r.$
@@
Recall that $2\lambda(A,\ell,row)$ and $2\lambda(B,k,col)$ are upper bounds on the maximum Hamming distance between a row of $A$ or a column of $B$ and its center in the $\ell$-center clustering or the $k$-center clustering computed by
$APPROXMMCLUS(A,B,\ell,k)$ in $PREPROCMMCLUS(A,B,\ell,k)$, respectively. Suppose $p\geq r.$ Step 1 takes $O(1)$ time. Since the $m$-th coordinate in the centers can be accessed in the matrix $A^{\prime}$ or $B^{\prime},$ respectively, computed by $APPROXMMCLUS(A,B,\ell,k)$, each of the two substeps in the block of the loop in Step 2 can be done in $O(1)$ time. Finally, since $|ind(A,j)|\leq 2\lambda(A,\ell,row)$, the block is iterated at most $2\lambda(A,\ell,row)$ times. Consequently, the whole Step 2 takes $O(\lambda(A,\ell,row))$ time. By symmetric arguments, if $p<r$ then Step 3 takes $O(\lambda(B,k,col))$ time. ∎
By putting Lemmata 3, 4, and 5 together, $(AB)^{\top}=B^{\top}A^{\top},$ and using the straightforward $O(sqt)$-time algorithm to multiply matrices of size $s\times q$ and $q\times t,$ we obtain our next main result.
@@
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. Given parameters $\ell\in[p]$ and $k\in[r],$ the matrices can be preprocessed by a simple deterministic algorithm in $O(pq\ell+pr)$ time if $p\geq r$ or $O(rqk+pr)$ time if $p\leq r$ such that a query asking for the exact value of a single entry $C_{ij}$ of the arithmetic matrix product $C$ of $A$ and $B$ can be answered in $O(\lambda(A,\ell,row))$ time if $p\geq r$ or $O(\lambda(B,k,col)))$ time if $p\leq r.$

### v2

@@ Abstract
We study applications of clustering (in particular the $k$-center clustering problem) in the design of efficient and practical deterministic algorithms for computing an approximate and the exact arithmetic matrix product of two 0-1 rectangular matrices $A$ and $B$ with clustered rows or columns, respectively. Let $\lambda_{A}$ and $\lambda_{B}$ denote the minimum maximum radius of a cluster in an $\ell$-center clustering of the rows of $A$ and in a $k$-center clustering of the columns of $B,$ respectively. In particular, when $A$ and $B$ are square matrices of size $n\times n$, we obtain the following results.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within an additive error of at most $2\lambda_{A}$ in $O(n^{2}\ell)$ time or at most $2\lambda_{B}$ in $O(n^{2}k)$ time.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(n^{2}\ell)$ time or $O(n^{2}k)$ time after which every query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time or $O(\lambda_{B})$ time, respectively.
@@ 1 Introduction
The arithmetic matrix product of two 0-1 matrices is closely related to the Boolean one of the corresponding Boolean matrices. For square $n\times n$ matrices, both can be computed in $O(n^{2.372})$ time [1, 23]. Both are basic tools in science and engineering. Unfortunately, no truly subcubic practical algorithms for any of them are known. Therefore, many researchers studied the complexity of these products for special input matrices, e.g., sparse or structured matrices [3, 4, 11, 15, 21, 24], providing faster and often more practical algorithms.
The method of multiplying matrices with clustered rows or columns, proposed for Boolean matrix product in [4] and subsequently generalized in [11, 15] and used in [2], relies on the construction of an approximate spanning tree of the rows of the first input matrix or the columns of the second input matrix in a Hamming space. Then, each column or each row of the product matrix is computed with the help of a traversal of the tree in time proportional to the total Hamming cost of the tree up to a logarithmic factor. Simply, the next entry in a column or a row in the product matrix can be obtained from the previous one in time roughly proportional to the Hamming distance between the consecutive (in the tree traversal) corresponding rows or columns of the first or the second input matrix, respectively. Thus, in case the entire tree cost is substantially subquadratic in $n,$ the total running time of this method becomes substantially subcubic provided that a good approximation of a minimum spanning tree of the rows of the first input matrix or the columns of the second one can be constructed in a substantially subcubic time. As for simplicity and practicality, a weak point of this method is that in order to construct such an approximation relatively quickly, it employs a randomized dimension reduction.
@@
Cohen and Lewis [8] and Drineas et al. [9] used random sampling to approximate arithmetic matrix product. Their papers provide an approximation $D$ of the matrix product $AB$ of two $n\times n$ matrices $A$ and $B$ such that $||AB-D||_{F}=O(||AB||_{F}/\sqrt{c})$, for a parameter $c>1$ (see also [21]). The approximation algorithm in [9] runs in $O(n^{2}c)$ time. Drineas et al. [9] also derived bounds on the entrywise differences between the exact matrix product and its approximation. Unfortunately, the best of these bounds is $\Omega(M^{2}n/\sqrt{c})$, where $M$ is the maximum value of an entry in $A$ and $B.$ By using a sketch technique, Sarlós [22] obtained the same Frobenius norm guarantees, also in $O(n^{2}c)$ time. However, he derived stronger individual upper bounds on the additive error of each entry $D_{ij}$ of the approximation matrix $D$. They are of the form $O(||A_{i*}||_{2}||B_{*j}||_{2}/\sqrt{c})$, where $A_{i*}$ and $B_{*j}$ stands for the $i$-th row of $A$ and the $j$-th column of $B,$ respectively, that hold with high probability. More recently, Pagh [21] presented a randomized approximation $\tilde{O}(n(n+c))$-time algorithm for the arithmetic product of $n\times n$ matrices $A$ and $B$ such that each entry of the approximate matrix product differs at most by $||AB||_{F}/\sqrt{c}$ from the correct one. His algorithm first compresses the matrix product to a product of two polynomials and then uses the fast Fourier transform to multiply the polynomials. Subsequently, Kutzkov [20] developed analogous deterministic algorithms employing different techniques. For approximation results related to sparse arithmetic matrix products, see [19, 21].
@@ 1.1 Our contributions
The $k$-center clustering problem in a Hamming space $\{0,1\}^{d}$ is for a set $P$ of $n$ points in $\{0,1\}^{d}$ to find a set $T$ of $k$ points in $\{0,1\}^{d}$ that minimize $\max_{v\in P}\min_{u\in T}\mathrm{ham}(v,u),$ where $\mathrm{ham}(v,u)$ stands for the Hamming distance between $v$ and $u,$ (i.e., the number of coordinate positions they differ from each other). Each center in $T$ induces a cluster consisting of all points in $P$ for which it is the nearest center.
@@
    A simple deterministic algorithm that approximates each entry of the arithmetic matrix product of $A$ and $B$ within an additive error of at most $2\lambda_{A}$ in $O(pq\ell+pr)$ time if $p\geq r$ or at most $2\lambda_{B}$ in $O(qrk+pr)$ time if $p\leq r$.
@@
    A simple deterministic preprocessing of the matrices $A$ and $B$ in $O(pq\ell+pr)$ time if $p\geq r$ or $O(qrk+pr)$ time if $p\leq r$ after which every query asking for the exact value of an arbitrary entry of the arithmetic matrix product of $A$ and $B$ can be answered in $O(\lambda_{A})$ time if $p\geq r$ or $O(\lambda_{B})$ time if $p\leq r.$
@@ 1.2 Techniques
All our main results rely on the classical, simple $2$-approximation algorithm for the $k$-center clustering problem (farthest-point clustering) due to Gonzalez [16] (see also Fact 1). Two of them rely also on the idea of updating the inner product of two vectors $a$ and $b$ in $\{0,1\}^{q}$ over the Boolean or an arithmetic semi-ring to that of two vectors $a^{\prime}$ and $b^{\prime}$ in $\{0,1\}^{q}$, where $a=a^{\prime}$ or $b=b^{\prime},$ in time roughly proportional to $\mathrm{ham}(a,a^{\prime})+\mathrm{ham}(b,b^{\prime})$. The idea has been used in [4, 11, 15]. As in the aforementioned papers, we combine it with a traversal of an approximate minimum spanning tree of the rows or columns of an input matrix in the Hamming space $\{0,1\}^{q}$, where $q$ is the length of the rows or columns (see also Lemma 6).
@@ 1.3 Paper organization
The next section contains basic definitions. Section 3 presents our approximation algorithm for the arithmetic product of two 0-1 matrices and the preprocessing enabling efficient answers to queries asking for the value of an arbitrary entry of the arithmetic product matrix. Section 4 is devoted to our algorithm for the exact arithmetic matrix product of two 0-1 matrices. We conclude with a short discussion on possible extensions of our results.
@@ 2 Preliminaries
The minimum-diameter $k$-clustering problem in a Hamming space $\{0,1\}^{d}$ is as follows: given a set $P$ of $n$ points in $\{0,1\}^{d}$, find a partition of $P$ into $k$ subsets $P_{1},P_{2},\dots,P_{k}$ that minimize $\max_{i\in[k]}\max_{v,u\in P_{i}}\mathrm{ham}(v,u).$ Note that the $k$-center clustering problem could be also termed as the minimum-radius $k$-clustering problem. It is known to be NP-hard and even NP-hard to approximate within $2-\epsilon$ for any constant $\epsilon>0$ [10, 13].
@@ 3 An approximate arithmetic matrix product of 0-1 matrices
procedure $APPROXMMCLUS(A,B,\ell)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$.
@@
Suppose a $2$-approximation algorithm for the $\ell$-center clustering is used in $APPROXMMCLUS(A,B,\ell)$ and $C$ stands for the arithmetic product of $A$ and $B.$ Then, for $1\leq i\leq p$ and $1\leq j\leq r,$ $|C_{ij}-D_{ij}|\leq 2\lambda(A,\ell,row)$.
@@
Recall that $p\geq r$ is assumed in the input to $APPROXMMCLUS(A,B,\ell)$. For $1\leq i\leq p$ and $1\leq j\leq r,$ $D_{ij}$ is the inner product of $cen_{\ell}(A_{i*})$, where $\mathrm{ham}(A_{i*},cen_{\ell}(A_{i*}))\leq 2\lambda(A,\ell,row)$, with $B_{*j}.$ Hence, $C_{ij}$, which is the inner product of $A_{i*}$ with $B_{*j},$ can differ at most by $2\lambda(A,\ell,row)$ from $D_{ij}.$
@@
$APPROXMMCLUS(A,B,\ell)$ can be implemented in
$O(pq\ell+pr+T(\ell,q,r))$ time.
@@
Recall that $p\geq r.$ Step 1, which includes the assignment of the closest center to each row of $A$, can be done in $O(pq\ell)$ time by using Fact 1, i.e., the classic algorithm of Gonzalez [16]. Step 2 takes $O(\ell q)$ time, which is $O(T(\ell,q,r))$ time. Finally, Step 3 takes $T(\ell,q,r)$ time while Step 4 can be done in $O(pr)$ time. Thus, the overall time is $O(pq\ell+pr+T(\ell,q,r)).$
We can use the straightforward $O(sqt)$-time algorithm for the multiplication of two matrices of sizes $s\times q$ and $q\times t,$ respectively. Since $T(\ell,q,r)=O(\ell qr)=O(\ell qp)$ if $p\geq r,$ Lemmata 1 and 2 yield the first part (1) of our first main result (Theorem 3.1 below), for $p\geq r.$ Its second part (2) for $p\leq r$ follows from the first part by $(AB)^{\top}=B^{\top}A^{\top}.$ Note that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Simply, we run $APPROXMMCLUS(B^{\top},A^{\top},k)$ in order to compute an approximation of the transpose of the arithmetic matrix product of $A$ and $B.$ Note also that a $k$-clustering of columns of $B$ is equivalent to a $k$-clustering of the rows of $B^{\top}$ and that $\lambda(B^{\top},k,row)=\lambda(B,k,col).$
@@
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. There is a simple deterministic algorithm which provides an approximation of all entries of the arithmetic matrix product of $A$ and $B$ within an additive error of at most:
@@
We slightly extend $APPROXMMCLUS(A,B,\ell)$ in order to obtain a preprocessing for answering queries about single entries of the arithmetic matrix product of $A$ and $B.$
procedure $PREPROCMMCLUS(A,B,\ell)$
Input: Two 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ and a positive integer $\ell$ not exceeding $p$.
Output: The $p\times r$ matrix $D$ returned by $APPROXMMCLUS(A,B,\ell)$, and for $1\leq i\leq p$, the set of coordinate indices $ind(A,i)$ on which $A_{i*}$ differs from its cluster center.
@@
    For $1\leq i\leq p,$ determine the set $ind(A,i)$ of coordinate indices on which $A_{i*}$ differs from $cen_{\ell}(A_{i*})$.
@@
$PREPROCMMCLUS(A,B,\ell)$ can be implemented in
$O(pq\ell+pr+T(\ell,q,r))$ time.
@@
Recall that $p\geq r.$ Step 1 can be done in $O(pq\ell+pr+T(\ell,q,r))$ time by Lemma 2. Step 2 can be easily implemented in $O(pq)$ time. ∎
@@
procedure $QUERYMMCLUS(A,B,\ell,i,j)$
Input: The preprocessing done by $PREPROCMMCLUS(A,B,\ell)$ for 0-1 matrices $A$ and $B$ of sizes $p\times q$ and $q\times r,$ respectively, where $p\geq r,$ $\ell\in[p]$, and two query indices $i\in[p]$ and $j\in[r].$
@@
$QUERYMMCLUS(A,B,\ell,i,j)$ is correct, i.e., the final value of $C_{ij}$ is the inner product of the $i$-th row $A_{i*}$ of $A$ and the $j$-th column $B_{*j}$ of $B.$
@@
$C_{ij}$ is initially set to $D_{ij}$, which is the inner product of the center assigned to $A_{i*}$ and $B_{*j}.$ Then, $C_{ij}$ is appropriately corrected by increasing or decreasing with $1$ for each coordinate index $m\in ind(A,i)$ which contributes $1$ to the inner product of $A_{i*}$ and $B_{*j}$ and 00 to the inner product of the center of $A_{i*}$ and $B_{*j}$ or vice versa. ∎
@@
$QUERYMMCLUS(A,B,\ell,k,i,j)$ takes $O(\lambda(A,\ell,row))$ time.
@@
Recall that $2\lambda(A,\ell,row)$ is an upper bound on the maximum Hamming distance between a row of $A$ $B$ and its center in the $\ell$-center clustering computed by $APPROXMMCLUS(A,B,\ell)$ in $PREPROCMMCLUS(A,B,\ell)$. Recall also that $p\geq r.$ Step 1 takes $O(1)$ time. Since the $m$-th coordinate in the centers can be accessed in the matrix $A^{\prime}$ computed by $APPROXMMCLUS(A,B,\ell)$, each of the two substeps in the block of the loop in Step 2 can be done in $O(1)$ time. Finally, since $|ind(A,j)|\leq 2\lambda(A,\ell,row)$, the block is iterated at most $2\lambda(A,\ell,row)$ times. Consequently, the whole Step 2 takes $O(\lambda(A,\ell,row))$ time. ∎
By putting Lemmata 3, 4, and 5 together, and using the straightforward $O(sqt)$-time algorithm to multiply matrices of size $s\times q$ and $q\times t,$ we obtain our next main result for $p\geq r.$ The case $p\leq r$ reduces to the case $p\geq r$ by $(AB)^{\top}=B^{\top}A^{\top}.$ Recall that then the number of rows in $B^{\top}$, which is $r$, is not less than the number of columns in $A^{\top}$, which is $p$. Also, we have $\lambda(B^{\top},k,row)=\lambda(B,k,col).$ We simply run $PREPROCMMCLUS(B^{\top},A^{\top},k)$ and
$QUERYMMCLUS(B^{\top},A^{\top},k,j,i)$ instead.
@@
Let $A$ and $B$ be two 0-1 matrices of sizes $p\times q$ and $q\times r,$ respectively. Given parameters $\ell\in[p]$ and $k\in[r],$ the matrices can be preprocessed by a simple deterministic algorithm in $O(pq\ell+pr)$ time if $p\geq r$ or $O(rqk+pr)$ time if $p\leq r$ such that a query asking for the exact value of a single entry $C_{ij}$ of the arithmetic matrix product $C$ of $A$ and $B$ can be answered in $O(\lambda(A,\ell,row))$ time if $p\geq r$ or $O(\lambda(B,k,col))$ time if $p\leq r.$